QUESTION 5

You are the owner of a business that manufactures notepads for conference venues. You also print branding details (venue name, address etc) on the notepads before delivery to the customer. You randomly select one hundred (100) orders and inspect each of the orders. Each order is considered a unit. You find the following defects:

- Two orders are only damaged (torn notepads)

- Three orders only have typos (spelling errors)

- One order is both damaged and also has typos

a) Calculate the defects per unit

Assume that the only way an order is considered defective is by being damaged or by having typos

b) How many defect opportunities are there?

c) If the process remains at this defect rate over the time it takes to produce one million orders, how many defects should you expect to find?

d) Suggest a single measure that you could take to prevent the defects described above. Base your answer on one of Deming’s 14 Points
Answer #1

(a) Defect per unit = total number of defects / total units

= (2+3+2) /100 = 0.07

(b) There are two defect opportunities, one having a typo and other being torn / damaged.

(c) Defect per million orders = defect per unit * 1000000

=0.07*1000000 = 70000

(d) Use of on the job training of employees is a measure that can be useful for making the employees engaged in packaging and typing work more apt, efficient and quality conscious. With continuous improvement oriented training, the company will be able to contain the defect rate.